OpenStudy (zmudz):

Find \(a/b\) when \(2\log{(a -2b)} = \log{a} + \log{b}\).

OpenStudy (michele_laino):

hint: I apply the properties of logarithm so I can write: \[{\left( {a - 2b} \right)^2} = ab\] so, after a simplification, we have: \[{a^2} + 4{b^2} - 4ab = ab\] or \[{a^2} - 5ab + 4{b^2} = 0\] now I divide both sides by b^2, so I get: \[{\left( {\frac{a}{b}} \right)^2} - 5\left( {\frac{a}{b}} \right) + 4 = 0\] then I make this variable change: \[z = \left( {\frac{a}{b}} \right)\] so I can rewrite the last equation as follows: \[{z^2} - 5z + 4\] please solve for z

OpenStudy (michele_laino):

oops.. \[{z^2} - 5z + 4 = 0\]